OpenStudy (anonymous):

what is the equation for the solution set (-5, 2)?

OpenStudy (anonymous):

a solution set with 2 values means that its a quadratic, or a polynomial possibly with repeated terms. indefinitely, the equation would be (x+5)(x-2) = x^2 + 3x - 10

OpenStudy (anonymous):

well x^2 + 3x - 10 = 0